OpenStudy (anonymous):

12/40 = n/60 what is N????

OpenStudy (anonymous):

\[\frac{ 12 }{ 40 }= \frac{ n }{ 60 }\] First you have to cross multiply 12*60=40n 720=40n Then Flip the equation. 40n=720 After that you Divide both sides by 40 40n/40=720/n N=18

OpenStudy (anonymous):

18 is correct